OpenStudy (anonymous):

Will give medal. got these ones wrong on the worksheet i turned in. Thank you :) 1. In how many different orders can 5 different appetizers of the 12 be brought to the table? 2. Assume that 3 customers arrive and order different appetizers to share from a choice of 12 appetizers. a. Does this problem involve permutations or combinations? b. Why? c. In how many possible ways can this be done? 3. Suppose that 5 customers arrive, and each orders a different appetizer to share from a choice of 12 appetizers. In how many ways can this be done?

OpenStudy (anonymous):

12*11*10*9*8 ways

OpenStudy (anonymous):

2)combinations

OpenStudy (anonymous):

3) order doesn't matter in combination

OpenStudy (anonymous):

4)12C3=12*11*10/3*2=

OpenStudy (anonymous):

5)12C5 ways

OpenStudy (anonymous):

@heyimrick

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

thank you so much. i gave u the medal, if u have any extra time i have 1 more question Suppose that 5 customers arrive, and each orders a different appetizer to share from a choice of 12 appetizers. In how many ways can this be done? @dg2

OpenStudy (anonymous):

12 c5 ways =12*10*11*9*8/2*3*4*5